OpenStudy (freeg13):

what is the slope of the line passing through the points (–1, 3) and (4 ,–7) A. 2 B. 3/4 C. -4/3 D. -2

OpenStudy (anonymous):

Hello, the slope of the line is given by: m = rise/run. You are given two points, can you calculate the rise and run between both points?

OpenStudy (anonymous):



OpenStudy (anonymous):



OpenStudy (anonymous):

@freeG13 come back please

OpenStudy (anonymous):

Welcome back @FreeG13. The rise is given by: -7 - 3 = -10. The run is given by: 4 - (-1) = 5. The slope is: rise/run: -10/5 = -2. Please let me know if any of this does not make sense. Learning is more important than the problem itself.

OpenStudy (freeg13):

thanks